1/1

Martha Müller-Grählert,

Mia hejmlando

translated by Manfredo Ratislavo

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Translation of the Low German poem “Mine Heimat” by Martha Müller-Grählert (∗1876-12-20 – †1939-11-18) into Esperanto by Manfredo Ratislavo (Manfred Retzlaff, Stettiner Str. 16, D-59302 Oelde, Germanio, ∗1938-11-04) on 1992-07.

MR-252-1 / Arg-375-737 (2013-01-08 09:18:56)

La melodio estas aŭdebla interalie tie: http: // www. youtube. com/ watch? v= gNmM_xEejEU kaj en http: // www. sjd-falken. de/ lieder/ wo_de_ostseewellen. shtml .

http://www.poezio.net/version?poem-id=375&version-id=,737